A glass of tea (175 g) cools from 85 ̊C down to comfortable room temperature 27. ̊C. How much energy does it release to the surroundings?
1
Expert's answer
2020-12-10T07:47:34-0500

Given that the mass of tea in the cup = 175 g

Final temperature = 27°С

Initial temperature = 85°С

Let us assume the specific heat of tea is equal to that of water = 4.184 J/g°C

Temperature change = Tfinal - Tinitial = (27 - 85) = - 58°C 

-Heat given out by tea = heat absorbed by the surroundings

Heat released into the surroundings =-mC(Tfinal - Tinitial) = -175 g × 4.184 J/g°C x (-55°C) = 40 271 J = 40.271 kJ